OpenStudy (anonymous):

ln(x) ---- 1/x Does this expression, if I take the limit for this expression as x --> 0+, does it give an indeterminate form? If so, how?

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0^{+}}x*ln(x)=0\]

OpenStudy (anonymous):

No I am asking if this gives 0/0, or infinity/infinity

OpenStudy (anonymous):

B/c then and only then can I use L'Hop. Rule

OpenStudy (anonymous):

yes is goes my friend, if you plug in 0 directtly you will get inifinty/inifinity

OpenStudy (anonymous):

How?

OpenStudy (anonymous):

ln(0) = - infinity, as x approaches from the right hand side, I understand that part..

OpenStudy (anonymous):

1/x, as x approaches 0 from the right hand side outputs infinity, OH OK. So you can put in infinity into the limit expresion? I thought I couldn't do that.

OpenStudy (anonymous):

1/0 is infinity, consider the graph: 

OpenStudy (anonymous):

well you thought wrong lol

OpenStudy (anonymous):

you want me wot finish that up for you or can you do it

OpenStudy (anonymous):

I can do it, but please PLEASE help me with this new question I am about to post. I am so close to getting the right answer, but I am still getting it "wrong" for some reason.

OpenStudy (anonymous):

okay, post it bro

OpenStudy (anonymous):

of course the graph i gave you was of 1/x

OpenStudy (anonymous):

you are solving \[\lim_{x\rightarrow 0^+} x\ln(x)\] yes? it is of the form \[0\times -\infty\]so by changing it to \[\lim_{x\rightarrow 0^+}\frac{\ln(x)}{\frac{1}{x}}\] you make it into \[\frac{-\infty}{\infty}\] and so you can use l'hopital

OpenStudy (anonymous):

Posted